The median age at first marriage for females increased from 24.5 years in 1995 to 25.1 years in 2000. Let 1995 be 5 and 2000 be year 10
a) Find the equation of the line through (5, 24.5) and (10, 25.1)
b) What do x and y represent in your equation?
c) Interpret the slope of this line
d) In what year will the median age be 30.
Your help is moslty appreciate, when I am in need. Thank you. Answer by stanbon(75887) 
 The median age at first marriage for females increased from 24.5 years in 1995 to 25.1 years in 2000. Let 1995 be 5 and 2000 be year 10
a) Find the equation of the line through (5, 24.5) and (10, 25.1)
slope = [25.1-24.5] / [10-5] = 0.6 / 5 = 3/25
------------------------
The form is y = mx+b where m=3/25 and y=24.5 when x=5
24.5 = (3/25)(5) +b
b = 239/10
----------
EQUATION:
y = (3/25)x + 239/10
-----------------------------
b) What do x and y represent in your equation?
x is # of years after 1990
y is median age at which women married
-------------------------
c) Interpret the slope of this line
For every 25 years after 1990 the median age at which wormen marry increases
by 3 years.
------------------------
d) In what year will the median age be 30.
30 = (3/25)x + (239/10)
(3/25)x = 6.1
x=50.83333
1990 + 50.8333.. = 2040.3333...
Rounding you get the year 2040
=================
Cheers,
Stan H.